What is the molar mass of 3.49 liters equal to? Which is the temperature of 285 K and the pressure of 207.7 kPa weigh 7.88 g?
1
Expert's answer
2019-01-04T03:18:11-0500

Answer:

PV = nRT

n = PV/RT

n = 207700*0.00349/8.31*285

n = 0.31 mole

M = m/n

M = 7.88/0.31 = 25.4 g/mole